# FreeBlend: Advancing Concept Blending with Staged Feedback-Driven Interpolation Diffusion

### v2

@@ 3.3 Staged Denoising Process
To achieve more flexible and precise control over image generation, motivated by Liew et al. (2022); Lin (2024); Ackermann & Li (2022), we divide the generation phase into three stages: the initialization stage, the blending stage, and the refinement stage. And we use two coefficients, $t_{s}$ and $t_{e}$, to denote the start and end timesteps of the blending stage, respectively.
At the initialization stage, the denoising process using Stable Diffusion begins with random noise $\epsilon\sim\mathcal{N}(0,I)$. In this stage, the blending latent is denoised based on the input unCLIP image embeddings, which define the basic layout of the subject and background in the image. Starting with $L^{(T)}_{b}=\epsilon$, the blending latent is iteratively updated according to the following:
$L^{(t-1)}_{b}=$ $\epsilon_{\theta}(L^{(t)}_{b},t,\varnothing)$ (6)
$+w\cdot\left[\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c}_{\text{stage}})-\epsilon_{\theta}(L_{b}^{(t)},t,\varnothing)\right],$
where $w$ denotes the classifier-free guidance (Ho & Salimans, 2022) scale, and $\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c})$ is the pre-trained U-Net model with conditional input $\mathbf{c}$ and timestep $t$.
At the blending stage, the initially formed latent is interpolated with auxiliary latents to incorporate original features. In this process, we employ a stepwise increasing interpolation strategy along with a feedback-driven mechanism to blend and filter out irrelevant features from the concepts.
At the refinement stage, the general appearance of $L_{b}$ has been completed. The focus now shifts to enhancing the finer details of the latent, making the result more natural and harmonious while eliminating any sense of disjointedness or artificial stitching. Using the same conditions as in the initialization stage, this phase continues to build upon the previous layout, further refining and improving it.
@@ 3.4 Stepwise Increasing Interpolation Strategy
In this section, we elaborate on how our method ensures semantic coherence. We interpolate two image latents with the current denoising latents to encode semantic information from both sources together, which shares similar intuition with MagicMix (Liew et al., 2022). The key challenge is to select appropriate blending ratio at each timestep. One issue of using a constant blending ratio is that we may result in blurring and unclear images, dramatically decreasing the generation quality. To address this, for a denoising process with $T$ timesteps, at timestep $t$, we apply a stepwise declining blending ratio $p$, which could be expressed as
@@
where $N$ is the number of concepts needed to blend, and the proportion of the $k$-th auxiliary latent $L^{(k)}_{a}$ is
$\lambda=\frac{1-p}{N}.$ (8)
Another challenge is the potential for biased or catastrophic forgetting of content. Specifically, the model should account for the varying influence of different concepts. For instance, when subject concepts such as “animals” are blended with background concepts like “plants”, the model might give more weight to the subject concepts, resulting in generated images with a higher representation of animals.
To address this, we introduce an additional hyper-parameter, denoted as $\gamma$, which regulates the contribution of each image embedding in the latent space. In the case of a blending task involving $N$ images, the interpolation process can be formulated as follows:
@@
Figure 3: The visualized results highlight the superiority of our approach over other methods. We provide a qualitative comparison of our method with three typical concept blending techniques: TEXTUAL (Melzi et al., 2023), Composable Diffusion (Liu et al., 2022) and MagicMix (Liew et al., 2022). For each method and pair, three images are presented. Compared to other methods, which suffer from rigid splicing, discordant compositions, or bias towards one concept, our method excels at seamlessly integrating the features of different concepts into a cohesive new object.
@@ 3.5 Feedback-Driven Mechanism
In terms of synthesis quality, setting the time proportion of the blending stage too high can lead to overlapping issues, particularly when the blending stage ends at a high value. This indicates that significant changes should not be introduced during the later stages of denoising, as this could cause unwanted fluctuations and overlapping effects. In previous methods (Liew et al., 2022), the blending stage ends earlier, which limits the ability for further blending and adjustment. To address this issue, we propose a feedback-driven mechanism.
@@
where $L^{(0,k)}_{a}$ is the embedding encoded from image.
Once $L^{(t)}_{b}$ is interpolated with $L^{(t,k)}_{a}$, the latter should be simultaneously updated using $L^{\prime(t)}_{b}$ as follows:
@@
$=\underbrace{(1-p)\cdot L_{a}^{(t,k)}}_{\text{inherits the original appearance}}$
$+\underbrace{p^{2}\cdot L^{(t)}_{b}}_{\text{integrates into the subject}}$
@@
$+\underbrace{p\cdot(1-p)\cdot\left[\frac{1}{N}\cdot\sum_{n=1}^{N}\gamma_{n}\cdot L_{a}^{(t,n)}\right]}_{\text{maintains balance}}.$
@@
After interpolation, $L_{a}$ should also be denoised as described in Equation 6, with $\mathbf{c}_{\text{stage}}$ being replaced by $\mathbf{c}_{\text{up}}$ or $\mathbf{c}_{\text{down}}$ according to the category of it.
@@
As shown in Equation 11, as the timesteps decrease, $p$ increases, which causes $L_{b}$ to take up a larger proportion of $L_{a}$. Consequently, in the later stages of blending, the interpolated $L_{a}$ increasingly resembles $L_{b}$, thus maintaining consistency over time. The average of all previous $L_{a}$ values contributes to the incorporation of general features, particularly during the middle stages of blending, due to the coefficient $p\cdot(1-p)$. Regarding $L^{(t,k)}_{a}$, it gradually declines in the final stages, thereby reducing the proportion of the original features. Through this approach, the auxiliary latents are updated by incorporating both the preceding auxiliary latents and the main blending latent.
@@
Table 1: Quantitative comparisons with other methods. Our method outperforms all other methods across all metrics, including main blending effect (CLIP-BS) , blending objects detection metric DINO-BS, image quality (CLIP-IQA), and better human preference (HPS).
@@
MagicMix (Liew et al., 2022) 8.3063$\pm$ 2.4757 0.2485$\pm$ 0.1575 0.4435$\pm$ 0.0980 0.2710$\pm$ 0.0290
Composable Diffusion (Liu et al., 2022) 6.1374$\pm$ 1.9454 0.2441$\pm$ 0.1655 0.4270$\pm$ 0.1082 0.2903$\pm$ 0.0277
unCLIP-T (Wang et al., 2023) 8.7433$\pm$ 3.1892 0.2214$\pm$ 0.1689 0.4436$\pm$ 0.1073 0.2384$\pm$ 0.0285
unCLIP-U (Wang et al., 2023) 8.7346$\pm$ 3.1577 0.2190$\pm$ 0.1703 0.4450$\pm$ 0.1061 0.2385$\pm$ 0.0280
TEXTUAL (Melzi et al., 2023) 7.8102$\pm$ 2.6852 0.2366$\pm$ 0.1537 0.4164$\pm$ 0.1143 0.2399$\pm$ 0.0230
UNET (Olearo et al., 2024) 8.5080$\pm$ 2.7110 0.2481$\pm$ 0.1628 0.3544$\pm$ 0.2785 0.2405$\pm$ 0.0326
FreeBlend(Ours) 9.1555$\pm$ 2.7134 0.2743$\pm$ 0.1586 0.5238$\pm$ 0.0975 0.2932$\pm$ 0.0316
@@
8.5577$\pm$ 2.6566
9.1555$\pm$ 2.7134
7.8102$\pm$ 2.6852
7.9888$\pm$ 2.5823
8.8170$\pm$ 2.7085
8.5080$\pm$ 2.7110
6.9255$\pm$ 2.4354
9.0260$\pm$ 2.7408
Table 3: Ablation study with different interpolation strategy. The results demonstrate that our increase strategy outperforms both the invariant and decline strategies.
Increase Invariant Decline CLIP-BS($\uparrow$) CLIP-IQA($\uparrow$)
9.1555$\pm$ 2.7134 0.5238$\pm$ 0.0975
7.8891$\pm$ 2.8732 0.4861$\pm$ 0.0990
8.5355$\pm$ 3.0644 0.4981$\pm$ 0.0836
Initialization Blending Refinement HPS($\uparrow$)
8.9552$\pm$ 2.7259
8.9832$\pm$ 2.8083
9.1313$\pm$ 2.6941
9.1555$\pm$ 2.7134
4 Experiments
4.1 Experimental Settings
Datasets Construction. We develop the Concept Text-Image Blending (CTIB) dataset, which includes 11,400 text-image pairs. These pairs originate from a collection of 380 text pairs distributed across 20 distinct categories, with each category containing 30 images. The categories are thoughtfully chosen to include a diverse array of real-world objects, highlighting the model’s capability to generate across various species and types.
Implementation Details. Our approach is executed with SD-2-$1$https://huggingface.co/stabilityai/stable-diffusion-2-1 for all baseline comparisons. For our suggested technique, we employ SD-2-1-unCLIP [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1-unclip, enabling image-based conditioning as an input. During the creation of image datasets, the prompt used is “a photo of a {class}”. To maintain a uniform resolution throughout the experiments, all images are resized to $768\times 768$ pixels. The experiments are conducted on a node equipped with 4 NVIDIA GeForce RTX A100 GPUs.
Metrics. Given the unique nature of our task compared to other generation tasks, we use four key metrics to evaluate the blending performance: (1) CLIP Blending Similarity (CLIP-BS) is the primary metric, which measures the distance and similarity between the blending results and the original concepts. (2) DINO Blending Similarity (DINO-BS) (Liu et al., 2025) assesses the detection score of blending objects or the combination of one object with features from another. (3) CLIP Image Quality Assessment (CLIP-IQA) (Wang et al., 2023) evaluates the image quality and the degree of match for the blending objects. (4) Human Preference Score (HPS) (Wu et al., 2023) measures human preferences for the blending results based on blending object prompts. Further explanation is provided in Section A.6.
4.2 Qualitative Comparisons
The results of our comparison with three existing typical concept blending algorithms are illustrated in Figure 3. We observe that using interpolated text embeddings tends to favor one category over the other, resulting in a failure to effectively integrate both elements into the final outcome. On the other hand, while Composable Diffusion successfully introduces both categories, it often leads to concepts co-occurring rather than achieving a harmonious blend. This is evident in cases like “car-cat”, where the concepts remain distinct rather than being seamlessly blended. MagicMix heavily depends on the shape similarity between the reference image and the corresponding category. As a result, it struggles when the categories are completely dissimilar, such as in the cases of “dog-neon light” and “orange-robot”. In these instances, the method fails to generate a meaningful blend, as the lack of visual similarity between the categories hinders the blending process. In contrast, our approach effectively merges both categories, producing well-integrated and balanced results. This enables smooth blending even when the categories are highly distinct, such as in “car-cat” or “dog-neon light”.
4.3 Quantitative Comparisons
We evaluate the performance on our CTIB dataset using four related metrics, with the results presented in Table 1. The results demonstrate that our method significantly outperforms all other approaches across all metrics. This indicates that the quality and the blending effect of our synthetic images achieved by our method is the most visually pleasing.
4.4 Ablation Study
Figure 4: Ablation study on the impact of $\gamma$. The results show that, in the first row, better blending is achieved on the left side, while the right side appears more spliced. In the second row, both sides exhibit a more visually appealing effect. The blending process, however, is inherently subjective, and users can adjust the parameter $\gamma$ to tailor the output according to their preferences. By modulating $\gamma$, users control the contribution of each concept, optimizing performance and addressing associated biases.
Figure 5: The results of user study across our method and three representative concept blending methods. The results demonstrate that our approach achieves the highest level of human preference.
Conditions for Denoising Process. Table 2 presents the results of our ablation study under different conditions. We classify the conditions into two main categories: text and image conditions. For text conditions, we use the TEXTUAL and UNET text prompt blending methods. For image conditions, we apply the unCLIP method along with either unCLIP-T (TEXTUAL-like) or unCLIP-U (UNET-like) image blending methods, where this approach treats the images as text conditions. The results demonstrate that the unCLIP-U method yields the best performance compared to other conditions, while text-based conditions tend to reduce the effectiveness of the generation.
Interpolation Strategy. Table 3 illustrates the impact of different interpolation strategies. In the blending stage, we vary the blending ratio between the interpolated auxiliary latents and the blending latent. From this table, we observe that the “increase” strategy yields a higher CLIP-BS and CLIP-IQA score. As $p$ increases, the blending latent gains more weight relative to the auxiliary latents, resulting in the final output exhibiting features of both kinds. However, the “invariant” method leads to a rigid adjustment, yielding the lowest performance. On the other hand, the “decline” strategy prioritizes the auxiliary latent space and avoids blending, which clearly results in ineffective blending.
Mitigation of bias. We introduce an additional parameter, $\gamma$, to reduce the bias toward specific categories in the generated output image. As shown in Figure 4, we vary $\gamma$ within the range $[0.5,1.5]$. Our observations show that as $\gamma$ moves from 0 to 2, the generated image gradually transitions from resembling reference 1 to favoring reference 2. This provides us with the flexibility to adjust the interpolation strength, allowing for precise control over the image’s characteristics and enabling the achievement of optimal results.
Staged Denoising Process. We validate the impact of our proposed staged denoising process in Table 4. The results demonstrate that both the initialization and refinement stages enhance the stability and quality of the generation. This highlights the effectiveness of the initialization stage in shaping the structure of the initial latents, while the refinement stage successfully adds finer details to the images.
4.5 User Study
We also conduct a user study to evaluate human preference on the blending results with images generated by different methods. Our survey comprised 6 sets of blending pairs. With the active participation of 50 volunteers, we successfully collected a total of 300 votes. The results, as presented in Figure 5, clearly demonstrate the dominance of our method. It received 192 votes, accounting for 64.00% of the total, far surpassing its competitors. Composable Diffusion obtained 63 votes (21.00%), MagicMix received 39 votes (13.00%), and Textual managed only 6 votes (2.00%). These findings not only affirm the effectiveness of our approach but also provide valuable insights into user preferences in the context of concept blending.
5 Conclusion
Blending multiple concepts into a single object based on diffusion models is interesting and valuable in 2D content creation but rather challenging, as it involves controlling the diffusion model to generate previously completely unseen objects. In this paper, we introduce FreeBlend, a training-free method to blend concepts with stark inherent semantic dissimilarities. At its core, FreeBlend employs the unCLIP model with images as conditions and utilizes a novel proposed interpolation strategy with feedback. Extensive qualitative and quantitative studies demonstrate that our method effectively blends disparate concepts, significantly surpassing the prior SoTA counterparts. Moving forward, exploring alternative ways to address the preference dynamics between different concepts may further improve the performance.
Impact Statement
This paper presents a contribution aimed at advancing the field of concept blending. We explicitly state that our work is intended for academic research and commercial applications, provided such uses are authorized by the author. However, we strongly emphasize that we do not condone, nor do we support, the use of our research to generate harmful, unethical, or unlawful content. The generation of unsafe or offensive images that violate ethical standards, legal regulations, or societal norms is strictly prohibited. As researchers, we believe it is our responsibility to ensure that our work is used for the betterment of society and in a manner that aligns with fundamental ethical principles.
Appendix A Appendix
A.1 T-SNE Analysis
In this section, we explore the CLIP embedding space of blended images. Given the high dimensionality of the embeddings, we apply t-SNE (Van der Maaten & Hinton, 2008) for dimensionality reduction and visualize the embeddings in 2D to illustrate their spatial relationships.
In Figure 7, it is evident that semantically related categories tend to be positioned closer to each other, while unrelated categories are spaced farther apart. In Figure 7, we observe that the blending images exhibit a strong relationship with the original classes. Notably, these blending images are generally biased towards one of the original classes, which reflects the inherent bias in the concept blending task. Specifically, we notice that the blending images are not located at the midpoint between the original classes. We hypothesize that this observation can be attributed to the following factors: a) t-SNE is a non-linear dimensionality reduction technique that aims to map similar points in high-dimensional space to nearby points in low-dimensional space, while preserving local structures rather than global relationships. Consequently, the position of points in the t-SNE plot may not necessarily correspond to the relative position of categories in the original feature space. b) The feature vector of the blended class may not be positioned directly between the two original classes. This could be due to: i) The blended class containing a mixture of features from both original classes, with the resulting feature vector being closer to one of the original classes in the feature space. ii) The blended feature may have a direction in the feature space that is relatively distant from both original classes. In certain cases, this could cause the blended class’s feature vector to extend further from the center, leading to its placement behind the original classes in the t-SNE plot, rather than being positioned between them.
A.2 Multiple Concepts Blending.
Our method is not limited to cases involving only two prompts. In Figure 8, we present examples where three or more prompts are incorporated. We utilize the TEXTUAL method to combine three prompts, as the UNET method is limited to handling only two prompts due to the constraints of the U-Net architecture. The blending effect is somewhat chaotic and abstract due to the significant differences and wide conceptual gap among the three prompts. Nevertheless, we managed to achieve relevant results and curated a selection of visually appealing outcomes to present.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
watermelon pineapple tree blending result epiphyllum violet rose blending result
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cat dog cactus blending result rainbow neon dog blending result
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
bee supercar yellow background blending result rainbow neon dog blending result
Figure 8: Multiple concepts blending results.
Figure 9: Ablation study of the feedback-driven mechanism. Removing the feedback module results in overlapping images, which disrupts the blending process and prevents successful integration.
@@ A.1 T-SNE Analysis
Figure 11: Adapting our method to style transfer task.
Figure 12: Screenshot of user study.
A.3 More Ablation Study
Feedback-Driven Mechanism. We conduct an ablation study in Table 5 to validate the effect of the feedback-driven mechanism. The results demonstrate the impact of our feedback-driven mechanism in blending different concepts. The DINO-BS score, however, is not optimal, which may be due to the surprising overlap of blended objects, thus boosting the score. We also present an intuitive visual comparison in Figure 9, where our effect is clearly significant.
@@ A.3 More Ablation Study
8.9211$\pm$ 3.0076 0.2860$\pm$ 0.1663 0.4874$\pm$ 0.1035 0.2866$\pm$ 0.0305
9.1555$\pm$ 2.7134 0.2743$\pm$ 0.1586 0.5238$\pm$ 0.0975 0.2932$\pm$ 0.0316
Visualization of Staged Denoising Process. Here, we provide a more detailed description of the staged denoising process employed in our method. To facilitate understanding, we visualize the intermediate latent representations, including the blending latent $L_{b}$ and the auxiliary latents $L_{a}^{(t,n)}$ ($n=0,1,2,...,N$) at various stages of the process. The results are presented in Figure 10. Specifically, we focus on timesteps 3 and 22, as they correspond to the initial and final stages where our method begins to exhibit its effects. These visualizations demonstrate how the latents evolve over time, gradually converging to a consistent structure. Through the feedback interpolation module, the blending latent and auxiliary latents seamlessly merge, ultimately resulting in a coherent output image.
A.4 Details of Dataset Construction

### v3

@@ 3.3 Staged Denoising Process
To enable more flexible and precise control over the image generation process, and inspired by previous works [liew2022magicmix, lin2024dreamsalon, ackermann2022high], we propose dividing the generation process into three stages: the initialization stage, the blending stage, and the refinement stage. We introduce two coefficients, $t_{s}$ and $t_{e}$, to represent the start and end timesteps of the blending stage, respectively.
At the initialization stage, the denoising process begins with random noise, $\epsilon\sim\mathcal{N}(0,I)$. In this stage, the blending latent is iteratively updated based on the input unCLIP image embeddings, which define the basic layout of the subject and background. The process starts with $L^{(T)}_{b}=\epsilon$, and the blending latent is updated according to the following equation:
$L^{(t-1)}_{b}=\epsilon_{\theta}(L^{(t)}_{b},t,\varnothing)+w\cdot\left[\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c}_{\text{stage}})-\epsilon_{\theta}(L_{b}^{(t)},t,\varnothing)\right],$ (3)
where $w$ denotes the classifier-free guidance [ho2021classifier] scale, and $\epsilon_{\theta}(L^{(t)}_{b},t,\mathbf{c})$ represents the pre-trained U-Net model with the conditional input $\mathbf{c}$ and timestep $t$.
In the blending stage, the focus shifts to incorporating original features into the initially formed latent. This process is guided by methods described in Section˜3.4 and Section˜3.5, which enable the effective blending and filtering of features, ensuring the integration of relevant information.
Finally, at the refinement stage, the general structure of the latent has been established, and the model’s focus moves toward enhancing finer details. The objective here is to improve the overall appearance, making the image more natural and cohesive, while addressing any disjointedness or artifacts from earlier stages. This phase continues to refine and improve the final image output, building on the layout established in the initialization stage.
@@ 3.4 Stepwise Increasing Interpolation Strategy
In this section, we elaborate on how our method ensures semantic coherence and consistency. We interpolate two image latents with the current denoising latents to encode semantic information from both sources together, which shares a similar intuition with MagicMix [liew2022magicmix]. The key challenge is to select an appropriate blending ratio at each timestep. One issue of using a constant blending ratio is that we may result in blurring and unclear images, dramatically decreasing the generation quality. To address this, for a denoising process with $T$ timesteps, at timestep $t$, we apply a stepwise declining blending ratio $p$, which could be expressed as
@@
and the proportion of the $k$-th auxiliary latent $L^{(k)}_{a}$ is
$\lambda=\frac{1-p}{N},$ (5)
where $N$ is the number of concepts needed to blend and we set it two. Another challenge is the potential for biased or catastrophic forgetting of content. Specifically, the model should account for the varying influence of different concepts. For instance, when subject concepts such as “animals” are blended with background concepts like “plants”, the model might give more weight to the subject concepts, resulting in generated images with a higher representation of animals.
To address this, we introduce an additional hyperparameter, denoted as $\gamma$, which regulates the contribution of each image embedding in the latent space. In the case of a blending task involving $N$ images, the interpolation process can be expressed as follows:
@@ 3.5 Feedback-Driven Mechanism
In terms of synthesis quality, setting the time proportion of the blending stage too high can lead to overlapping issues, particularly when the blending stage is extended too long. This indicates that the interpolation methods used in previous approaches are unable to handle significant changes, resulting in incomplete processing and causing overlapping effects. In contrast, earlier methods [liew2022magicmix] terminate the blending stage prematurely, which restricts further blending and fine-tuning. To address this, we propose a feedback-driven mechanism that strikes an optimal balance between these two extremes.
@@
where $L^{(0,k)}_{a}$ is the embedding encoded from image. Once $L^{(t)}_{b}$ is interpolated with $L^{(t,k)}_{a}$, the latter should simultaneously undergo a feedback update using $L^{\prime(t)}_{b}$ as follows:
@@
$=\underbrace{(1-p)\cdot L_{a}^{(t,k)}}_{\text{inherits the original appearance}}+\underbrace{p^{2}\cdot L^{(t)}_{b}}_{\text{integrates into the subject}}+\underbrace{p\cdot(1-p)\cdot\left[\frac{1}{N}\cdot\sum_{n=1}^{N}\gamma_{n}\cdot L_{a}^{(t,n)}\right]}_{\text{maintains balance}}.$
@@
After interpolation, $L_{a}$ should also be denoised normally, with $\mathbf{c}_{\text{stage}}$ being replaced by $\mathbf{c}_{\text{up}}$ or $\mathbf{c}_{\text{down}}$ according to the category of it.
@@
As shown in Equation˜8, as the timesteps decrease, $p$ increases, which causes $L_{b}$ to take up a larger proportion of $L_{a}$. Consequently, in the later stages of blending, the interpolated $L_{a}$ increasingly resembles $L_{b}$, thus maintaining consistency over time. The average of all previous $L_{a}$ values contributes to the incorporation of general features, particularly during the middle stages of blending, due to the coefficient $p\cdot(1-p)$. Regarding $L^{(t,k)}_{a}$, it gradually declines in the final stages, thereby reducing the proportion of the original features. Through this approach, the auxiliary latents are updated by incorporating both the preceding auxiliary latents and the main blending latent.
4 Experimental Results
4.1 Experimental Settings
Datasets Construction. To ensure a fair comparison between text and image-conditioned methods while minimizing the deviation caused by external datasets, we create the Concept Text-Image Reference (CTIR) dataset. This dataset consists of 20 categories, each containing 30 images. The categories have been carefully selected to represent a wide variety of real-world objects, showcasing the model’s ability to generate content across different species and types. For evaluation purposes, we develop the Concept Text-Image Blending (CTIB) dataset, which includes 5,700 text-image pairs. These pairs are drawn from a set of 190 text prompts, distributed across the 20 categories, with each category containing 30 images.
Implementation Details. We use SD-2-1111https://huggingface.co/stabilityai/stable-diffusion-2-1 as the backbone model for the baseline comparisons. For the pipeline, we employ SD-2-1-unCLIP [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1-unclip, which enables image-based conditioning as an input. To maintain a uniform resolution throughout the experiments, all images are resized to $768\times 768$ pixels. The experiments are conducted on a node equipped with 4 NVIDIA GeForce RTX A100 GPUs.
Metrics. Given the unique nature of our task compared to other generation tasks, we use four key metrics to evaluate the blending performance: (1) CLIP Blending Similarity (CLIP-BS) is the primary metric, which measures the CLIP [radford2021learning] distance and similarity between the blending results and the original concepts. (2) DINO Blending Similarity (DINO-BS) [liu2025grounding] assesses the detection score of blending objects or the combination of one object with features from another. (3) CLIP Image Quality Assessment (CLIP-IQA) [wang2023exploring] evaluates the image quality and the degree of match for the blending objects. (4) Human Preference Score (HPS) [wu2023human] measures human preferences for the blending results based on blending object prompts. Further explanation is provided in Appendix˜G.
Figure 3: At the top are the concepts, and on the left are the methods we compare. Each row shows the results, with three images per method and concept pair, evaluating our method against five blending methods. Unlike others, which can suffer from rigid splicing, discordant compositions, or concept bias, our method smoothly integrates features from different concepts into a cohesive new object.
4.2 Qualitative Comparisons
The qualitative comparison of our method with five existing approaches applicable to the concept blending task is shown in Figure˜3. We observe that interpolated text embeddings [melzi2023does] tend to favor one category, failing to integrate both concepts effectively. Composable Diffusion [liu2022compositional] introduces both categories but often results in them co-occurring rather than blending harmoniously. MagicMix [liew2022magicmix] heavily relies on shape similarity between the reference image and the categories, struggling when the categories are dissimilar, such as “dog-neon light” or “orange-robot”. unCLIP-T and unCLIP-U [ramesh2022hierarchical] blend the two concepts well, but their blending direction and extent are suboptimal, failing to generate a meaningful blend when the categories lack visual similarity. In contrast, our method effectively merges both categories, producing well-integrated results even with highly distinct concepts like “car-cat” or “dog-neon light”.
4.3 Quantitative Comparisons
We evaluate our method on our designed CTIB dataset with four key metrics, with the results presented in Table˜1. The results demonstrate that our method significantly outperforms all other approaches across all metrics. This indicates that the quality and the blending effect of our synthetic images achieved by our method is the most visually pleasing.
@@ 4.3 Quantitative Comparisons
Table 1: Quantitative comparisons with other methods. Our method outperforms all other methods across all metrics, including the main blending effect (CLIP-BS), the blending objects detection metric DINO-BS, image quality (CLIP-IQA), and better human preference (HPS).
@@
[…]
Table 2: Ablation study with different interpolation strategies. The results demonstrate that our increase strategy outperforms both the invariant and decline strategies.
Increase Invariant Decline CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
[…]
Initialization Blending Refinement CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
[…]
Figure 4: Ablation study on the impact of $\gamma$. The results show that, in the first row, better blending is achieved on the left side, while the right side appears more spliced. In the second row, both sides exhibit a more visually appealing effect. The blending process, however, is inherently subjective, and users can adjust the parameter $\gamma$ to tailor the output according to their preferences. By adjusting $\gamma$, users can control the contribution of each concept, thereby mitigating associated biases.
4.4 Ablation Study
Mitigation of bias. Different concepts inherently exhibit biases in Stable Diffusion, likely due to the training datasets and their relevance to specific concepts or content. So we introduce an additional parameter, $\gamma$, to reduce the bias toward specific categories in the generated output image. As shown in Figure˜4, we vary $\gamma$ within the range $[0.5,1.5]$. Our observations show that as $\gamma$ moves from 0 to 2, the generated image gradually transitions from resembling reference 1 to favoring reference 2. This provides us with the flexibility to adjust the interpolation strength, allowing for precise control over the image’s characteristics and enabling the achievement of optimal results.
Interpolation Strategy. Table˜2 illustrates the impact of different interpolation strategies. In the blending stage, we vary the blending ratio between the interpolated auxiliary latents and the blending latent. From this table, we observe that the “increase” strategy yields a higher CLIP-BS and CLIP-IQA score. As $p$ increases, the blending latent gains more weight relative to the auxiliary latents, resulting in the final output exhibiting features of both kinds. However, the “invariant” method leads to a rigid adjustment, yielding the lowest performance. On the other hand, the “decline” strategy prioritizes the auxiliary latent space and avoids blending, which clearly results in ineffective blending.
Staged Denoising Process. We validate the impact of our staged denoising process in Table˜3. The results demonstrate that both the initialization and refinement stages enhance the stability and quality of the generation. This highlights the effectiveness of the initialization stage in shaping the structure of the initial latents, while the refinement stage successfully adds finer details to the images.
Figure 5: Ablation study of the feedback mechanism: removing the feedback module causes image overlap, disrupting blending and preventing integration.
Feedback-Driven Mechanism. We conduct an ablation study in Table˜7 to validate the effect of the feedback-driven mechanism. The results demonstrate the impact of our feedback mechanism in blending different concepts. The DINO-BS score, however, is not optimal, which may be due to the surprising overlap of blended objects, thus boosting the score. We also present an intuitive visual comparison in Figure˜5, where our effect is clearly significant.
4.5 User Study
We also conduct a user study to evaluate human preference for the blending results with images generated by different methods. Our survey comprises 50 sets of blending pairs. With the active participation of 26 volunteers, we successfully collected a total of 1300 votes. The results are displayed in Table˜4. These findings not only affirm the effectiveness of our approach but also provide valuable insights into user preferences in the context of concept blending.
Table 4: User study results comparing FreeBlend with three other methods. Higher values indicate better user preference.
Method MagicMix TEXTUAL Composable Diffusion FreeBlend(Ours) Total
User Votes ($\uparrow$) 49(3.77%) 42(3.23%) 127(9.77%) 1082(83.2%) 1300
5 Conclusion
Blending multiple concepts into a single object based on diffusion models is both interesting and valuable for 2D content creation. However, it is quite challenging, as it requires controlling the diffusion model to generate objects that have never been seen before. In this paper, we introduce FreeBlend, a training-free method to blend concepts with stark inherent semantic dissimilarities. At its core, FreeBlend employs the unCLIP model with images as conditions and utilizes a novel proposed interpolation strategy with feedback. Extensive qualitative and quantitative studies demonstrate that our method effectively blends disparate concepts, significantly surpassing the prior SoTA counterparts. Moving forward, exploring alternative ways to address the preference dynamics between different concepts may further improve the performance.
Appendix A T-SNE Analysis
In this section, we explore the CLIP embedding space of blended images. Given the high dimensionality of the embeddings, we apply t-SNE [van2008visualizing] for dimensionality reduction and visualize the embeddings in 2D to illustrate their spatial relationships with the corresponding concepts.
In Figure˜7, it is evident that semantically related categories tend to be positioned closer to each other, while unrelated categories are spaced farther apart. This is reasonable given the contrastive training objective of CLIP. In Figure˜7, we observe that the blending images exhibit a strong relationship with the original classes. Notably, these blending images are generally biased towards one of the original classes, which reflects the inherent bias in the concept blending task.
Specifically, we notice that the blending images are not located at the midpoint between the original classes. We hypothesize that this observation can be attributed to the following factors: a) t-SNE is a non-linear dimensionality reduction technique that aims to map similar points in high-dimensional space to nearby points in low-dimensional space, while preserving local structures rather than global relationships. Consequently, the position of points in the t-SNE plot may not necessarily correspond to the relative position of categories in the original feature space. b) The feature vector of the blended class may not be positioned directly between the two original classes. This could be due to: i) The blended class containing a mixture of features from original classes, with the resulting feature vector being closer to one of the original classes in the feature space. ii) The blended feature may have a direction in the feature space that is relatively distant from both original classes. In certain cases, this could cause the blended class’s feature vector to extend further from the center, leading to its placement behind the original classes in the t-SNE plot, rather than being positioned between them.
Appendix B Comparison on TIF Dataset
To provide a convincing comparison with other image-conditioned methods, we conduct experiments on the TIF dataset [xiong2024novel]. Table˜5 demonstrates that our method consistently achieves superior results.
Table 5: Quantitative comparisons with other image-conditioned methods on TIF dataset.
[…]
Appendix C More Ablation Study
Conditions for Denoising Process. Table˜6 presents the results of our ablation study under different conditions. For text conditions, we use the TEXTUAL and UNET text prompt blending methods for comparison. For image conditions, we apply the unCLIP method along with either unCLIP-T (TEXTUAL-like) or unCLIP-U (UNET-like) image blending methods. We also explore the method of simultaneously using text and images. The results demonstrate that the unCLIP-U method yields the best performance compared to other conditions, while text-based conditions tend to reduce the effectiveness of the generation.
@@ Appendix C More Ablation Study
[…]
@@
Visualization of Staged Denoising Process. Here, we provide a more detailed description of the staged denoising process employed in our method. To facilitate understanding, we visualize the intermediate latent representations, including the blending latent $L_{b}$ and the auxiliary latents $L_{a}^{(t,n)}$ (n=0,1,2,,Nn=0,1,2,...,N) at various stages of the process. The results are presented in Figure˜8. Specifically, we focus on timesteps 3 and 22, as they correspond to the initial and final stages where our method begins to exhibit its effects. These visualizations demonstrate how the latents evolve over timesteps, gradually converging to a consistent structure. Through the feedback interpolation module, the blending latent and auxiliary latents seamlessly merge, ultimately resulting in a coherent output image.
@@
[…]
Appendix D Details of Dataset Construction